WIDE-RANGE VISIBLE SPECTROSCOPY OF HIGH-SPIN ALKALI TRIMERS ON HELIUM DROPLETS, AND INTERPRETATION OF THEIR ELECTRONIC STRUCTURE

JOHANN NAGL, ANDREAS W. HAUSER, GERALD AUBÖCK, OLIVIER ALLARD, \underline{CARLO CALLEGARI}, and WOLFGANG E. ERNST, Institute of Experimental Physics, TU Graz, Petersgasse 16, 8010 Graz, Austria/EU.

We use helium droplets to assemble all possible combinations of trimers of potassium and rubidium (K_{3}, K_{2}Rb, KRb_{2}, Rb_{3}) in a quartet-spin state, and we measure their electronic excitation spectra in the range 10600–17400 cm^{−1}. We use two methods to separate overlapping spectra: visible-visible V-type double-resonance and mass-selective beam depletion. We observe and assign the laser-induced-fluorescence spectra of ten new bands. Assignment is backed by high-level ab initio calculations of the electronic levels structure of these molecules. Both experiment and calculations show highly regular patterns which we elucidate in terms of an electron-shell structure similar in spirit to that observed in Na clusters.^{a}

aW. D. Knight, K. Clemenger, W. A. de Heer, W. A. Saunders, M. Y. Chou, and M. L. Cohen Phys. Rev. Lett. 52, 2141 (1984).